Given r1=3mi+4mj.dertemine the magnitude and direction angles of the position vector

1
Expert's answer
2022-05-25T10:27:04-0400
"\vec r=(3\hat i+4\hat j)\:\rm m"

The magnitude

"r=\sqrt{3^2+4^2}=5\;\rm m"

The direction angles

"\alpha=\cos^{-1}\frac{3}{5}=53^\circ"

"\beta=\cos^{-1}\frac{4}{5}=37^\circ"